A $100,000 bond payable is issued on June 1, 20X1 at 106.

The bond comes due in exactly 5 years.

The bond pays annual cash interest of 12% with payments every June 1 and December 1.

Assuming the straight-line method is being used, how much (rounded) should be reported for the liability as of December 31, 20X1?
Answer #1

Solution:

Issue price of bond = $100,000*106%= $106,000

Premium on issue of bond = $106,000 - $100,000 = $6,000

Premium amortization in 20X1 = $6,000/5*7/12 = $700

Bond liability to be reported as of December 31, 20X1 = $106,000 - $700 = $105,300

Interest payable to be reported on December 31, 20X1 = $100,000*12%*1/12 = $1,000